OpenStudy (anonymous):

Please Help! Parameterize the plane that contains the three points (3,−5,−2), (10,−12,−10), and (25,0,5). r(s,t)=

OpenStudy (anonymous):

I first assigned the points to P Q and R and then found PQ and PR and then put P+u<PQ>+v<PR> and I got <3+7s+22t,-5-7s+5t,-2-8s+7s> but got it wrong!

jimthompson5910 (jim_thompson5910):

what did you get for vector PQ

jimthompson5910 (jim_thompson5910):

btw, this page may help you out http://math.harvard.edu/~ytzeng/worksheet/0926_sol.pdf see page 3

OpenStudy (anonymous):

I got <7,-7,-8>

OpenStudy (anonymous):

Yes i looked at that and Im stll not sure what i dd wrong

jimthompson5910 (jim_thompson5910):

how about PR?

OpenStudy (anonymous):

22,5,7

jimthompson5910 (jim_thompson5910):

ok, so far, so good

jimthompson5910 (jim_thompson5910):

I'm assuming you tried <3+7s+22t,-5-7s+5t,-2-8s+7t>and not <3+7s+22t,-5-7s+5t,-2-8s+7s> right?

OpenStudy (anonymous):

i did it twice and got it wrong twice

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

does it specify at all which variables are the free variables (like x and y or x and z)?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

hmm well I managed to find a calculator that will give the equation of the plane

jimthompson5910 (jim_thompson5910):

that equation is -x-25y+21z-80 = 0 -x-25y+21z= 80 solve for x: -x-25y+21z= 80 -x = 80+25y-21z x = -80-25y+21z

jimthompson5910 (jim_thompson5910):

so if you let y = s and z = t, then you will get <x, y, z> = <-80-25y+21z, y, z> <x, y, z> = <-80-25s+21t, s, t>

jimthompson5910 (jim_thompson5910):

unfortunately there are many ways to parameterize the plane

OpenStudy (anonymous):

can you please provide me with an example

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

as in please show me a way o parameterize this

jimthompson5910 (jim_thompson5910):

well you could take the cross product of vectors PQ and PR

jimthompson5910 (jim_thompson5910):

that will give you the normal vector to the plane

OpenStudy (anonymous):

what would i do then

jimthompson5910 (jim_thompson5910):

what do you get when you do the cross product

OpenStudy (anonymous):

gimme a sec

OpenStudy (anonymous):

-89i+113j+189k

jimthompson5910 (jim_thompson5910):

somewhat close, but that's not correct

OpenStudy (anonymous):

what is it then?

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

ok